OpenStudy (lena772):

Check my answer? The number of subscribers y to a newspaper after t years is shown by the equation below. y = 75(0.95)^t Which conclusion is correct about the number of subscribers to the newspaper? It increased by 75% every year. It decreased by 75% every year. It increased by 5% every year. It decreased by 5% every year. It's D right?

OpenStudy (kittiwitti1):

Hmmmm..... what subject is this?

OpenStudy (lena772):

It's supposed to be y=75(0.95)^t

OpenStudy (lena772):

It's Algebra

OpenStudy (kittiwitti1):

Ah, so it's \[y=75(0.95)^t\]then?

OpenStudy (lena772):

yes

OpenStudy (lena772):

I'm pretty sure it's D i just want to make sure

OpenStudy (kainui):

This should be explicitly asked, @Lena772 is it: y = 75(0.95)t or y = 75(0.95)^t

OpenStudy (lena772):

I said it was 75(0.95)^t

OpenStudy (kittiwitti1):

Not in the question, no you didn't.

OpenStudy (lena772):

... ok.

OpenStudy (agent0smith):

You are right, lena.

OpenStudy (kainui):

Then yes, you're right @Lena772 Since if you plug in 0 for t, it's 1 for the (.95) part, so it'll be just itself. plug in 1 and you get that it's now .95 of the original, which is smaller. plug in 2 and you get .95 of what you plugged in for 1, since it's twice. So you can see that it's really decreasing by 5% each time.

OpenStudy (lena772):

Thanks @agent0smith

OpenStudy (lena772):

Thanks @Kainui and @kittiwitti1

OpenStudy (kittiwitti1):

So you might want to correct that.... and according to my graphing calculator it's right.

OpenStudy (kittiwitti1):

I'm down here lol /waves

OpenStudy (agent0smith):

You don't need to graph it, or plug in any numbers The "base" (it has the exponent attached), which is 0.95 in this case, subtract 1 from it to find out the decimal growth. Then convert to a percentage.

OpenStudy (agent0smith):

eg \[\Large P = 1000(0.08)^t\]what would be the growth rate? or \[\Large A = 20(1.09)^t\]what about that one?